### v1

@@ 5.1.4 Thermonuclear X-ray bursts
The sudden intense X-ray emission from a thermonuclear burst will interact with the surrounding accretion environment [for a review, see 238], more so if the flux or fluence are higher such as in superbursts or intermediate-duration bursts [239, 240, 241, 242]. In contrast, the accretion flow will affect the burst radiation. This interaction is expressed through multiple effects. Bursts can cool the corona (causing a hard X-ray deficit, 243), alter the accretion flow (leading to an increase of the accretion spectrum, 244 or a decrease 245, e.g.,), and can produce reflection features[246] as the burst photons bounce off the accretion disk. Although these interaction mechanisms likely coexist during thermonuclear bursts, their relative dominance remains unclear and is likely to depend on system-specific conditions. The hard X-ray deficit has been detected in 10 bursters to date, out of 120 known, often in systems with low accretion rates. Furthermore, the observed saturation of the deficit fraction at 50% in several bursters implies the coexistence of two distinct hard X-ray emission mechanisms [247]: a quasi-spherical corona undergoing burst-induced cooling and a persistent jet structure near the polar caps that remains unaffected during bursts. Accretion spectrum enhancement (the ‘ effect’, [244, 248]) is nearly ubiquitous in bright bursts (e.g., for peak luminosities higher than 10% of the Eddington limit). However, the correlation between burst-induced accretion spectrum enhancement and burst emission varies significantly between burst sources [249], indicating a source-dependent coupling mechanism between these phenomena, likely depending on the geometry and perspective of the accretion flow. There exists a degeneracy among the interaction mechanisms described above. For example, both the increased accretion rate model and the disk reflection scenario could explain the observed enhancement in the soft X-ray band [250]. The accretion circumstance can also have an impact on the burst emission. For instance, for the most luminous bursts with photospheric expansion (PRE), there is a peculiarity that the bolometric flux seems to keep increasing through the Eddington-limited phase. This is observed in bursts at a high mass accretion rate, but not for bursts with a faint persistent emission, which has been predicted theoretically but only observed in a single burst [251]. If this effect is due to disk obscuration, the phenomenon indicates that the anisotropy of the burst emission is accretion-rate dependent, which could be evidence of the truncated disk in the low/hard state, and the phenomenon could be used to estimate the inner disk radius. An enhancement in hard X-ray emission has been observed during short-duration bursts in two bursts. This phenomenon is thought to result from photon upscattering by either the accretion disk corona or the boundary layer between the accretion disk and neutron star. Notably, burst sources that exhibit elevated persistent luminosities, such as Cyg X-2 and Cir X-1, frequently demonstrate atypical characteristics, including the absence of spectral softening. These deviations from standard burst behavior are similarly thought to originate from coronal-scattering processes, where the persistent emission’s hot electron population interacts with the burst photons. With its unmatched effective area below a few keV, the SFA will uniquely probe dominant burst-accretion interactions and probe the accretion physics for burst sources with relatively low interstellar absorption. The SFA observations of bursts will enable a more accurate determination of the enhanced accretion rate and disk reflection parameters, providing a more detailed description of the accretion conditions throughout the bursts.
@@ 5.1.5 Compact and transient jets in BHXRBs
Theoretical models and numerical simulations suggest that the accretion flows surrounding black holes generate the magnetic fields and energy necessary to launch relativistic jets [e.g. 252, 253, 254, 255]. In transient BHXRBs, two types of jet are observed: a compact jet during the hard-state transition and a transient jet during the hard-to-soft-state transition. Significant progress has been achieved through multi-wavelength observations, offering valuable insights into jet activities and their couplings with accretion processes [see reviews such as 256]. Nevertheless, a comprehensive understanding of the mechanisms of jet formation and acceleration, as well as its large-scale propagation, is still lacking. Likewise, the physical processes that govern individual jet components and their connection with the rapid changes of accretion flow continue to present unresolved questions in both stellar-mass and super-massive accreting BHs.
@@
However, high-cadence observations of eXTP will enable us to capture X-ray spectral, timing, and polarimetric variations associated with changes of the corona, while radio imaging will track the onset and propagation of downstream transient jet components. The prolonged exposure capacity of eXTP is particularly well suited to capture the precise moment of the launch of the transient jet during the state transition [e.g. 257]. When combined with multi-epoch radio interferometry observations, eXTP promises to unravel the causal relationship between corona evolution and transient jet ejection. This synergy will advance our understanding of the physical process underlying state transitions, such as geometric and/or magnetic reconfiguration of the corona, and energy drainage into jet launching.
@@ 5.2.1 Gamma-ray binaries
Gamma-ray binaries are a rare subclass of high-mass X-ray binaries (HMXBs) comprising a compact object (neutron star or black hole) and a massive O- or Be-type companion star. These systems derive their name from their characteristic spectral energy distributions (SEDs), which peak at high energies (HE; MeV) and extend to very high energies (VHE; GeV) [258]. To date, only nine confirmed gamma-ray binaries have been identified within the Milky Way, including one system discovered in the Large Magellanic Cloud [259]. Their broadband emission exhibits strong orbital modulation, attributed to relativistic particle acceleration in colliding stellar winds or jets, synchrotron radiation at lower energies (radio to X-rays), and inverse Compton scattering (ICS) of stellar photons at the HE and VHE bands [see, e.g., 258, and references therein].
@@
In the 0.5-10 keV energy band, gamma-ray binaries typically show nonthermal spectra characterized by hard power-law emission with photon indices in the range , with no measured Fe line component and flux modulation over orbital timescales [258]. The usual assumption is that this X-ray emission is attributed to synchrotron radiation from relativistic electrons accelerated within the system, possibly in the shock formed by the interaction of a pulsar wind and the wind of the massive companion star, or in jets. Other models have been proposed where the X-ray emission is due to inverse Compton emission [266, 267, 259].
@@
Finally, we note that the recent literature reports few marginal detections of transient gamma-ray sources spatially coinciding with accreting pulsars in HMXBs that are neither microquasars nor colliding-wind binaries [277, 278, 279]. These outliers might represent only a small fraction of the total population of HMXBs, which, on the basis of X-ray observations, is known to be significantly larger than those of conventional gamma-ray binaries. Although numerous hypotheses have been proposed, the mechanism driving gamma-ray emission from accreting X-ray pulsars remains poorly understood [280, and references therein]. In this context, the superb sensitivity, excellent timing resolution, and advanced spectral capabilities of the SFA and PFA will be critical for obtaining deeper insights into the accretion flow dynamics and magnetic field geometry of these sources.
@@ 5.2.2 Supergiant High Mass X-ray Binaries
X-rays originating near compact objects are subject to absorption by surrounding matter, followed by subsequent re-emission. This phenomenon, known as “X-ray reprocessing”, manifests primarily through characteristic fluorescence lines, such as Fe Kα at 6.4 keV, Fe XXV He triplets around 6.7 keV, and FeXXVI Lya line at 6.97 keV. These features are related to the distribution of the accreting matter, which could therefore serve as a powerful diagnostic probe for the environment in binary systems [295, 296, 297]. eXTP, with a large effective area of SFA, is quite suitable for studying line emissions, especially in faint sources and low states of variable sources (see the comparison between eXTP/SFA and XMM-Newton/PN in Figure 14). For example, Cen X-3 shows enhanced Fe lines in its low states, suggesting that the suppressed flux is due to obscuration by dense matter rather than a reduction of the accretion rate [e.g. 298]. In addition, in a scenario where primary X-ray emission is heavily obscured and the observed flux is dominated by reprocessing, a high degree of polarization is expected, which may reach 25% as suggested by IXPE in thick AGNs Cyg X-3 and Compton [271, 299]. In this case, PFA polarization studies will play a pivotal role in understanding the structure and geometry of accretion flows.
@@ 6 Ultraluminous X-ray sources
Intermediate-mass black holes (IMBHs) have been proposed as possible accretors in some ULXs. The large effective area and superior timing resolution of the SFAs enable detailed spectral and timing analyses, offering a pathway to test this scenario. In particular, QPOs serve as a potential diagnostic tool. Although QPOs have been well established in stellar-mass black hole binaries (BHBs; ), they have also been detected in AGN hosting supermassive black holes (–), such as RE J1034396 [316] and 1ES 1927654 [317], albeit at much lower frequencies due to the larger black hole masses. Using SFAs’ enhanced timing capabilities, ULXs can be probed for QPOs on a wide range of frequencies, allowing mass-scale comparisons with BHBs and AGNs. Such analyses could provide critical evidence for the presence of IMBHs in ULX systems.
@@
PFA sensitivity and good telemetry constraints will enable detailed studies of both highly-obscured (high inclination) sources and extreme super-Eddington transients. Polarization constraints will refine funnel opening angles, enabling proper comparison between different sources and enabling direct constraints on the mass accretion rate estimates. Furthermore, these constraints will help distinguish between funnel-like (high polarization degree orthogonal to the jet) versus corona-like geometries (polarization degree smaller than about 10% and direction along the jet), which represents an important diagnostic for the possibility of testing the strong gravity effects (see White Paper on Strong Gravity).
@@ 7.1 Radio-Quiet AGNs
Active galactic nuclei (AGN) are powerful X-ray emitters located at the center of the host galaxy and fueled by accretion onto a supermassive black hole (SMBH). According to the standard model of AGN, the accreting material forms an optically thick and geometrically thin accretion disk [318] that emits in the optical / UV waveband. The detection of significant amounts of X-ray flux released as a power law and extending up to hundreds of keV, requires the presence of a corona of hot (with energy up to keV, e.g. 327) and optically thin plasma capable of inverse-Compton scattering the optical/UV seed photons received from the disk into the X-ray band [e.g. 328, 329, 330]. A fraction of the primary X-ray radiation is scattered back towards the disk, where it interacts with the relatively colder accreting gas producing a reflection spectrum [e.g. 331, 332]. The fraction of hard X-ray flux that irradiates the disk depends on the exact coronal geometry, which is currently unknown. One of the strongest features in the X-ray reflection spectrum is the fluorescence Fe K line. The profile of this spectral line is highly distorted by relativistic effects if the emission is from the innermost radii of an accretion disk extending close to the last stable orbit [e.g. 333]. The reflection spectrum and the relativistic Fe K line are, therefore, powerful diagnostics of the strong gravity field close to the SMBH. One of the defining properties of AGN is their extreme X-ray variability. The timescales of the fastest registered variations suggest that the hot corona is very compact and is located in the innermost regions of the accretion flow.
@@
Parameters such as accretion rate and SMBH mass provide critical insights into the coupling mechanisms of matter and energy between the disk and coronae in the vicinity of SMBHs [e.g., 334, 335, 336, 337, 338].
The outer regions of AGN are populated by clouds of gas, the so-called broad line region (BLR, e.g. 339), moving in roughly Keplerian motion around the BH and producing broad optical/UV emission lines. The BLR is delimited by an equatorial, dusty, and probably clumpy pc-scale torus [e.g. 340, 341]. All these components are enveloped by a narrow line region (NLR) where the clouds orbit at lower speeds. Although the bulk of their emission is at longer wavelengths, these outer regions may contribute to the X-ray spectrum with reflection features due to reprocessing, and with absorption from ionized and cold material.
The large effective area and the CCD-class spectral resolution of the SFA onboard eXTP coupled with the complementary simultaneous high-quality X-ray polarization data from the PFA will allow significant steps forward in our understanding of the AGN structure. The unprecedented SFA spectral sensitivity will allow hundreds of AGN to be observed down to a flux of in the local universe with a very good accuracy (i.e., a signal-to-noise ratio in excess of 100). In particular, eXTP capabilities will be fundamental to answer the following questions.
@@
Below a soft excess above the extrapolation of the hard X-ray power-law continuum is ubiquitously observed in unobscured type 1 AGN, with an occurrence rate [e.g. 342]. This feature is particularly prominent in narrow-line Seyfert 1 (NLSy1) galaxies, which are also characterized by high Eddington ratios and strong X-ray variability. Despite its prevalence, the physical origin of soft excess remains an open question. The proposed “warm corona” models suggest that the soft excess represents the hard tail of a broad-band (from optical/UV to soft X-rays) continuum component, produced via Compton up-scattering of the disk photons in a warm ( keV), optically thick () plasma [e.g., 343, 344, 345], possibly covering the inner disk. On the other hand, reflection models explain the soft excess as relativistic blurred ionized reflection [e.g., 331, 346, 332]. Recent theoretical [e.g., 347, 348] and observational studies [e.g., 349] suggest that both components may coexist in AGN. However, the exact contribution of each component to the soft X-ray excess and the way in which they vary as a function of the physical parameters of the AGN (e.g., BH mass, and accretion rate) remain poorly understood. Disentangling these two components is crucial for understanding the origin of the soft excess, precisely constraining the geometry of the innermost accretion flow, determining the BH spin, and providing us with a more comprehensive understanding of the energy dissipation processes in the vicinity of the BH. In addition, resolving the nature of the soft excess will be key to deciphering the physics behind the extreme properties of some sources, such as NLSy1 galaxies.
With its unprecedented effective area, the eXTP SFA offers a unique opportunity to address this problem. The left panel of Figure 16 compares the simulated X-ray spectrum of
an AGN resembling the NLSy1 galaxy Mrk 335, with a soft X-ray flux of , observed over 100 ks with XMM-Newton EPIC-pn and eXTP SFA. Both simulated spectra are obtained by assuming the same hybrid model. This model consists of a linear combination of emission from the warm corona (compTT) and ionized reflection (relxilllp). While the SFA spectrum clearly exhibits a substantial improvement in signal-to-noise, the middle and right panels of Figure 16 further illustrate the significant enhancement in parameter constraints enabled by the SFA. These panels present confidence contours for key parameters of the ionized reflection component (i.e. the reflection strength and the ionization parameter ) as well as for the warm corona component (i.e. the soft X-ray, keV, warm corona-to-ionized reflection flux ratio and the warm corona temperature ). With an effective area four times larger than that of XMM-Newton EPIC-pn, eXTP SFA reduces parameter uncertainties by approximately half, demonstrating its superior capability for constraining AGN X-ray emission mechanisms.
@@
In addition, the reflected emission is expected to be intrinsically polarized up to [e.g. 352, 353, 354]. Under certain conditions (relatively high inclination and/or high reflection fraction), the reflection component may contribute a significant fraction of the measured polarization degree in the keV.
For instance, some NLSy1 galaxies exhibit sharp drops in X-ray flux that coincide with the emergence of complex spectral features, which have been attributed to a dominant reflection component [e.g. 355].
Detailed spectro-polarimetric analysis with simultaneous eXTP PFA and SFA data will allow the contribution of reflection to be more precisely estimated. For example, spectropolarimetric simulations based on a simple toy model (including the same spectral components as those shown in Figure 16, but incorporating polarization in the reflection component using KYNSTOKES [352, 356], while assuming that the hot / warm corona emission is not polarized) suggest that simultaneous eXTP PFA and SFA observations could further reduce statistical uncertainty in the reflection strength by 10–20%.
@@
The dominant contribution to the X-ray spectrum of radio-quiet AGN originates from the hot corona, whose physical nature is currently unknown. The proposed models mainly invoke magnetic processes, suggesting that the corona might be a compact region powered by magnetic reconnection events [e.g. 357], the base of a (failed) jet [358], or a slab-like structure sandwiching the accretion disk. Each of these models predicts distinct coronal geometries, making it crucial to obtain strong observational constraints on its structure to uncover its physical origin.
Constraints on coronal geometry can be obtained through various methods. However, previous attempts based on X-ray spectral and timing analysis have encountered significant degeneracies [e.g. 332], which limits the ability to derive robust conclusions about the corona’s size, shape, and location relative to the accretion disk. Although larger effective area detectors, such as the eXTP SFA, and advances in spectral-timing models offer promising improvements, overcoming these degeneracies may remain a challenge. A breakthrough in this regard has been the advent of X-ray polarimetric analysis with IXPE [65]. Since polarization properties are highly sensitive to the geometry of the emitting region [e.g. 354], measuring the polarization degree and angle of X-ray emission in radio-quiet AGN provides a powerful tool to break the degeneracies inherent in spectral and timing analyses alone. By combining polarization data with traditional X-ray observations, it is possible to place much stronger constraints on the structure and orientation of the corona. However, current efforts to constrain the geometry of the corona in type 1 radio-quiet AGN through polarimetric analysis of IXPE data have faced several challenges. One major limitation is IXPE’s relatively small effective area, which has allowed for significant constraints on a few of the brightest Seyfert galaxies in the local universe [e.g. 359]. Additionally, achieving robust constraints on the different spectral components requires a strong synergy with other X-ray telescopes to obtain simultaneous high-quality X-ray spectra.
The advent of eXTP represents a major step forward in spectro-polarimetric analysis. With an effective area approximately four times that of IXPE, eXTP PFA will significantly enhance the sensitivity and precision of polarimetric measurements, reducing uncertainties and allowing the detection of polarization signals in weaker sources. Furthermore, the simultaneous observations provided by eXTP PFA and SFA will enable comprehensive spectro-polarimetric fits for a broad range of AGN, eliminating the need to rely on external instruments for complementary spectral information. This unprecedented capability will offer a more complete and self-consistent picture of the coronal geometry and emission mechanisms in AGNs.

### v2

@@ 5.1.4 Thermonuclear X-ray bursts
The sudden intense X-ray emission from a thermonuclear burst will interact with the surrounding accretion environment [for a review, see 239], more so if the flux or fluence are higher such as in superbursts or intermediate-duration bursts [240, 241, 242, 243]. In contrast, the accretion flow will affect the burst radiation. This interaction is expressed through multiple effects. Bursts can cool the corona (causing a hard X-ray deficit, 244), alter the accretion flow (leading to an increase of the accretion spectrum, 245 or a decrease 246, e.g.,), and can produce reflection features[247] as the burst photons bounce off the accretion disk. Although these interaction mechanisms likely coexist during thermonuclear bursts, their relative dominance remains unclear and is likely to depend on system-specific conditions. The hard X-ray deficit has been detected in 10 bursters to date, out of 120 known, often in systems with low accretion rates. Furthermore, the observed saturation of the deficit fraction at 50% in several bursters implies the coexistence of two distinct hard X-ray emission mechanisms [248]: a quasi-spherical corona undergoing burst-induced cooling and a persistent jet structure near the polar caps that remains unaffected during bursts. Accretion spectrum enhancement (the ‘ effect’, [245, 249]) is nearly ubiquitous in bright bursts (e.g., for peak luminosities higher than 10% of the Eddington limit). However, the correlation between burst-induced accretion spectrum enhancement and burst emission varies significantly between burst sources [250], indicating a source-dependent coupling mechanism between these phenomena, likely depending on the geometry and perspective of the accretion flow. There exists a degeneracy among the interaction mechanisms described above. For example, both the increased accretion rate model and the disk reflection scenario could explain the observed enhancement in the soft X-ray band [251]. The accretion circumstance can also have an impact on the burst emission. For instance, for the most luminous bursts with photospheric expansion (PRE), there is a peculiarity that the bolometric flux seems to keep increasing through the Eddington-limited phase. This is observed in bursts at a high mass accretion rate, but not for bursts with a faint persistent emission, which has been predicted theoretically but only observed in a single burst [252]. If this effect is due to disk obscuration, the phenomenon indicates that the anisotropy of the burst emission is accretion-rate dependent, which could be evidence of the truncated disk in the low/hard state, and the phenomenon could be used to estimate the inner disk radius. An enhancement in hard X-ray emission has been observed during short-duration bursts in two bursts. This phenomenon is thought to result from photon upscattering by either the accretion disk corona or the boundary layer between the accretion disk and neutron star. Notably, burst sources that exhibit elevated persistent luminosities, such as Cyg X-2 and Cir X-1, frequently demonstrate atypical characteristics, including the absence of spectral softening. These deviations from standard burst behavior are similarly thought to originate from coronal-scattering processes, where the persistent emission’s hot electron population interacts with the burst photons.
With its unmatched effective area below a few keV, the SFA will uniquely probe dominant burst-accretion interactions and probe the accretion physics for burst sources with relatively low interstellar absorption. The SFA observations of bursts will enable a more accurate determination of the enhanced accretion rate and disk reflection parameters, providing a more detailed description of the accretion conditions throughout the bursts.
@@ 5.1.5 Compact and transient jets in BHXRBs
Theoretical models and numerical simulations suggest that the accretion flows surrounding black holes generate the magnetic fields and energy necessary to launch relativistic jets [e.g. 253, 254, 255, 256]. In transient BHXRBs, two types of jet are observed: a compact jet during the hard-state transition and a transient jet during the hard-to-soft-state transition. Significant progress has been achieved through multi-wavelength observations, offering valuable insights into jet activities and their couplings with accretion processes [see reviews such as 257]. Nevertheless, a comprehensive understanding of the mechanisms of jet formation and acceleration, as well as its large-scale propagation, is still lacking. Likewise, the physical processes that govern individual jet components and their connection with the rapid changes of accretion flow continue to present unresolved questions in both stellar-mass and super-massive accreting BHs.
@@
High-cadence observations of eXTP will enable us to capture X-ray spectral, timing, and polarimetric variations associated with changes of the corona, while radio imaging will track the onset and propagation of downstream transient jet components. The prolonged exposure capacity of eXTP is particularly well suited to capture the precise moment of the launch of the transient jet during the state transition [e.g. 258]. When combined with multi-epoch radio interferometry observations, eXTP promises to unravel the causal relationship between corona evolution and transient jet ejection. This synergy will advance our understanding of the physical process underlying state transitions, such as geometric and/or magnetic reconfiguration of the corona, and energy drainage into jet launching.
@@ 5.2.1 Gamma-ray binaries
Gamma-ray binaries are a rare subclass of high-mass X-ray binaries (HMXBs) comprising a compact object (neutron star or black hole) and a massive O- or Be-type companion star. These systems derive their name from their characteristic spectral energy distributions (SEDs), which peak at high energies (HE; MeV) and extend to very high energies (VHE; GeV) [259]. To date, only 7 confirmed gamma-ray binaries have been identified within the Milky Way, with one additional system discovered in the Large Magellanic Cloud [260]. Their broadband emission exhibits strong orbital modulation, attributed to relativistic particle acceleration in colliding stellar winds or jets, synchrotron radiation at lower energies (radio to X-rays), and ICS of stellar photons at the HE and VHE bands [see, e.g., 259, and references therein].
@@
In the 0.5–10 keV energy band, gamma-ray binaries typically show nonthermal spectra characterized by hard power-law emission with photon indices in the range , with no measured Fe line component and flux modulation over orbital timescales [259]. The usual assumption is that this X-ray emission is attributed to synchrotron radiation from relativistic electrons accelerated within the system, possibly in the shock formed by the interaction of a pulsar wind and the wind of the massive companion star, or in jets. Other models have been proposed where the X-ray emission is due to inverse Compton emission [267, 268, 260].
@@
Finally, we note that the recent literature reports few marginal detections of transient gamma-ray sources spatially coinciding with accreting pulsars in HMXBs that are neither microquasars nor colliding-wind binaries [278, 279, 280]. These outliers might represent only a small fraction of the total population of HMXBs, which, on the basis of X-ray observations, is known to be significantly larger than those of conventional gamma-ray binaries. Although numerous hypotheses have been proposed, the mechanism driving gamma-ray emission from accreting X-ray pulsars remains poorly understood [281, and references therein]. In this context, the superb photon collecting area, excellent timing resolution, and good spectral capabilities of the SFA and PFA will be critical for obtaining deeper insights into the accretion flow dynamics and magnetic field geometry of these sources.
@@ 5.2.2 Supergiant High Mass X-ray Binaries
X-rays originating near compact objects are subject to absorption by surrounding matter, followed by subsequent re-emission. This phenomenon, known as “X-ray reprocessing”, manifests primarily through characteristic fluorescence lines, such as Fe Kα at 6.4 keV, Fe XXV He triplets around 6.7 keV, and FeXXVI Lya line at 6.97 keV. These features are related to the distribution of the accreting matter, which could therefore serve as a powerful diagnostic probe for the environment in binary systems [296, 297, 298]. eXTP, with a large effective area of the SFA, is quite suitable for studying line emissions, especially in faint sources and low states of variable sources (see the comparison between eXTP-SFA and XMM-Newton-PN in Figure 14). For example, Cen X-3 shows enhanced Fe lines in its low states, suggesting that the suppressed flux is due to obscuration by dense matter rather than a reduction of the accretion rate [e.g. 299]. In addition, in a scenario where primary X-ray emission is heavily obscured and the observed flux is dominated by reprocessing, a high degree of polarization is expected, which may reach 25% as suggested by IXPE in thick AGNs Cyg X-3 and Compton [272, 300]. In this case, PFA polarization studies will play a pivotal role in understanding the structure and geometry of accretion flows.
@@ 6 Ultraluminous X-ray sources
Intermediate-mass black holes (IMBHs) have been proposed as possible accretors in some ULXs. The large effective area and superior timing resolution of the SFA enable detailed spectral and timing analyses, offering a pathway to test this scenario. In particular, QPOs serve as a potential diagnostic tool. Although QPOs have been well established in stellar-mass black hole binaries (BHBs; ), they have also been detected in AGN hosting supermassive black holes (–), such as RE J1034396 [317] and 1ES 1927654 [318], albeit at much lower frequencies due to the larger black hole masses. Using SFA’ enhanced timing capabilities, ULXs can be probed for QPOs on a wide range of frequencies, allowing mass-scale comparisons with BHBs and AGNs. Such analyses could provide critical evidence for the presence of IMBHs in ULX systems.
@@
The large effective area and good telemetry constraints of the eXTP-PFA will enable detailed studies of both highly-obscured (high inclination) sources and extreme super-Eddington transients. Polarization constraints will refine funnel opening angles, enabling proper comparison between different sources and enabling direct constraints on the mass accretion rate estimates. Furthermore, these constraints will help distinguish between funnel-like (high polarization degree orthogonal to the jet) versus corona-like geometries (polarization degree smaller than about 10% and direction along the jet), which represents an important diagnostic for the possibility of testing the strong gravity effects (see [4]).
@@ 7.1 Radio-Quiet AGNs
Active galactic nuclei (AGN) are powerful X-ray emitters located at the center of the host galaxy and fueled by accretion onto a supermassive black hole (SMBH). According to the standard model of AGN, the accreting material forms an optically thick and geometrically thin accretion disk [319] that emits in the optical / UV waveband. The detection of significant amounts of X-ray flux released as a power law and extending up to hundreds of keV, requires the presence of a corona of hot (with energy up to keV, e.g. 328) and optically thin plasma capable of ICS the optical/UV seed photons received from the disk into the X-ray band [e.g. 329, 330, 331]. A fraction of the primary X-ray radiation is scattered back towards the disk, where it interacts with the relatively colder accreting gas producing a reflection spectrum [e.g. 332, 333]. The fraction of hard X-ray flux that irradiates the disk depends on the exact coronal geometry, which is currently unknown. Recently studies have shown that there are non-thermal electrons in the AGN hot corona [334, 335, 336], and could lead to gamma-ray emissions [337]. A much larger extended corona is proposed to explain the observed gamma-ray emission. One of the strongest features in the X-ray reflection spectrum is the fluorescence Fe K line. The profile of this spectral line is highly distorted by relativistic effects if the emission is from the innermost radii of an accretion disk extending close to the last stable orbit [e.g. 338]. The reflection spectrum and the relativistic Fe K line are, therefore, powerful diagnostics of the strong gravity field close to the SMBH. One of the defining properties of AGN is their extreme X-ray variability. The timescales of the fastest registered variations suggest that the hot corona is very compact and is located in the innermost regions of the accretion flow.
@@
Parameters such as accretion rate and SMBH mass provide critical insights into the coupling mechanisms of matter and energy between the disk and coronae in the vicinity of SMBHs [e.g., 339, 340, 341, 342, 343].
The outer regions of AGN are populated by clouds of gas, the so-called broad line region (BLR, e.g. 344), moving in roughly Keplerian motion around the BH and producing broad optical/UV emission lines. The BLR is delimited by an equatorial, dusty, and probably clumpy pc-scale torus [e.g. 345, 346]. All these components are enveloped by a narrow line region (NLR) where the clouds orbit at lower speeds. Although the bulk of their emission is at longer wavelengths, these outer regions may contribute to the X-ray spectrum with reflection features due to reprocessing, and with absorption from ionized and cold material.
The large effective area and the CCD-class spectral resolution of the SFA onboard eXTP coupled with the complementary simultaneous high-quality X-ray polarization data from the PFA will allow significant steps forward in our understanding of the AGN structure. The unprecedented effective area of the SFA will allow hundreds of AGN to be observed down to a flux of in the local universe with a very good accuracy (i.e., a signal-to-noise ratio in excess of 100). In particular, eXTP capabilities will be fundamental to answer the following questions.
@@
Below a soft excess above the extrapolation of the hard X-ray power-law continuum is ubiquitously observed in unobscured type 1 AGN, with an occurrence rate [e.g. 347]. This feature is particularly prominent in narrow-line Seyfert 1 (NLSy1) galaxies, which are also characterized by high Eddington ratios and strong X-ray variability. Despite its prevalence, the physical origin of soft excess remains an open question. The proposed “warm corona” models suggest that the soft excess represents the hard tail of a broad-band (from optical/UV to soft X-rays) continuum component, produced via Compton up-scattering of the disk photons in a warm ( keV), optically thick () plasma [e.g., 348, 349, 350], possibly covering the inner disk. On the other hand, reflection models explain the soft excess as relativistic blurred ionized reflection [e.g., 332, 351, 333]. Recent theoretical [e.g., 352, 353] and observational studies [e.g., 354] suggest that both components may coexist in AGN. However, the exact contribution of each component to the soft X-ray excess and the way in which they vary as a function of the physical parameters of the AGN (e.g., BH mass, and accretion rate) remain poorly understood. Disentangling these two components is crucial for understanding the origin of the soft excess, precisely constraining the geometry of the innermost accretion flow, determining the BH spin, and providing us with a more comprehensive understanding of the energy dissipation processes in the vicinity of the BH. In addition, resolving the nature of the soft excess will be key to deciphering the physics behind the extreme properties of some sources, such as NLSy1 galaxies.
With its unprecedented effective area, the eXTP-SFA offers a unique opportunity to address this problem. The left panel of Figure 16 compares the simulated X-ray spectrum of
an AGN resembling the NLSy1 galaxy Mrk 335, with a soft X-ray flux of , observed over 100 ks with XMM-Newton EPIC-pn and eXTP SFA. Both simulated spectra are obtained by assuming the same hybrid model. This model consists of a linear combination of emission from the warm corona (compTT) and ionized reflection (relxilllp). While the SFA spectrum clearly exhibits a substantial improvement in signal-to-noise, the middle and right panels of Figure 16 further illustrate the significant enhancement in parameter constraints enabled by the SFA. These panels present confidence contours for key parameters of the ionized reflection component (i.e. the reflection strength and the ionization parameter ) as well as for the warm corona component (i.e. the soft X-ray, keV, warm corona-to-ionized reflection flux ratio and the warm corona temperature ). With an effective area four times larger than that of XMM-Newton EPIC-pn, eXTP-SFA reduces parameter uncertainties by approximately half, demonstrating its superior capability for constraining AGN X-ray emission mechanisms.
@@
In addition, the reflected emission is expected to be intrinsically polarized up to [e.g. 357, 358, 359]. Under certain conditions (relatively high inclination and/or high reflection fraction), the reflection component may contribute a significant fraction of the measured polarization degree in the keV.
For instance, some NLSy1 galaxies exhibit sharp drops in X-ray flux that coincide with the emergence of complex spectral features, which have been attributed to a dominant reflection component [e.g. 360].
Detailed spectro-polarimetric analysis with simultaneous eXTP-PFA and SFA data will allow the contribution of reflection to be more precisely estimated. For example, spectropolarimetric simulations based on a simple toy model (including the same spectral components as those shown in Figure 16, but incorporating polarization in the reflection component using KYNSTOKES [357, 361], while assuming that the hot / warm corona emission is not polarized) suggest that simultaneous eXTP-PFA and SFA observations could further reduce statistical uncertainty in the reflection strength by 10–20%.
@@
The dominant contribution to the X-ray spectrum of radio-quiet AGN originates from the hot corona, whose physical nature is currently unknown. The proposed models mainly invoke magnetic processes, suggesting that the corona might be a compact region powered by magnetic reconnection events [e.g. 362], the base of a (failed) jet [363], or a slab-like structure sandwiching the accretion disk. Each of these models predicts distinct coronal geometries, making it crucial to obtain strong observational constraints on its structure to uncover its physical origin.
Constraints on coronal geometry can be obtained through various methods. However, previous attempts based on X-ray spectral and timing analysis have encountered significant degeneracies [e.g. 333], which limits the ability to derive robust conclusions about the corona’s size, shape, and location relative to the accretion disk. Although larger effective area detectors, such as the eXTP-SFA, and advances in spectral-timing models offer promising improvements, overcoming these degeneracies may remain a challenge. A breakthrough in this regard has been the advent of X-ray polarimetric analysis with IXPE [65]. Since polarization properties are highly sensitive to the geometry of the emitting region [e.g. 359], measuring the polarization degree and angle of X-ray emission in radio-quiet AGN provides a powerful tool to break the degeneracies inherent in spectral and timing analyses alone. By combining polarization data with traditional X-ray observations, it is possible to place much stronger constraints on the structure and orientation of the corona. However, current efforts to constrain the geometry of the corona in type 1 radio-quiet AGN through polarimetric analysis of IXPE data have faced several challenges. One major limitation is IXPE’s relatively small effective area, which has allowed for significant constraints on a few of the brightest Seyfert galaxies in the local universe [e.g. 364]. Additionally, achieving robust constraints on the different spectral components requires a strong synergy with other X-ray telescopes to obtain simultaneous high-quality X-ray spectra.
The advent of eXTP represents a major step forward in spectro-polarimetric analysis. With an effective area approximately five times that of IXPE, eXTP-PFA will significantly enhance the sensitivity and precision of polarimetric measurements, reducing uncertainties and allowing the detection of polarization signals in weaker sources. Furthermore, the simultaneous observations provided by eXTP-PFA and SFA will enable comprehensive spectro-polarimetric fits for a broad range of AGN, eliminating the need to rely on external instruments for complementary spectral information. This unprecedented capability will offer a more complete and self-consistent picture of the coronal geometry and emission mechanisms in AGNs.